# Stepwise Reasoning Error Disruption Attack of LLMs

### v3

@@ Abstract
Large language models (LLMs) have made remarkable strides in complex reasoning tasks, but their safety and robustness in reasoning processes remain unexplored, particularly in third-party platforms that facilitate user interactions via APIs. Existing attacks on LLM reasoning are constrained by specific settings or lack of imperceptibility, limiting their feasibility and generalizability. To address these challenges, we propose the Stepwise rEasoning Error Disruption (SEED) attack, which subtly injects errors into prior reasoning steps to mislead the model into producing incorrect subsequent reasoning and final answers. Unlike previous methods, SEED is compatible with zero-shot and few-shot settings, maintains the natural reasoning flow, and ensures covert execution without modifying the instruction. Extensive experiments on four datasets across four different models demonstrate SEED’s effectiveness, revealing the vulnerabilities of LLMs to disruptions in reasoning processes. These findings underscore the need for greater attention to the robustness of LLM reasoning to ensure safety in practical applications. Our code is available at: https://anonymous.4open.science/r/Reasoning-Attack-4663/
@@
Stepwise Reasoning Error Disruption Attack of LLMs
Jingyu Peng‡§11footnotemark: 1, Maolin Wang§11footnotemark: 1, Xiangyu Zhao§ 22footnotemark: 2 , Kai Zhang, Wanyu Wang§, Pengyue Jia§, Qidong Liu§, Ruocheng Guo, Qi Liu 22footnotemark: 2 University of Science and Technology of China, § City University of Hong Kong, Independent Researcher jypeng28@mail.ustc.edu.cn morin.wang@my.cityu.edu.hk
@@ 1 Introduction
Large language models (LLMs) have remarkably improved complex tasks by adopting various enhanced reasoning approaches Besta et al. (2024); Yang et al. (2024); Yao et al. (2024). These approaches have boosted their performance and drawn attention to the trustworthiness of the reasoning processes, including faithfulness Lanham et al. (2023); Turpin et al. (2024), fairness Shaikh et al. (2023), and safety Xu et al. (2024).
@@
Exploiting such vulnerability in LLMs introduces two fundamental challenges: feasibility and imperceptibility. Technically, unlike traditional adversarial attack methods, which often leverage internal information of target models such as gradients and logits, state-of-the-art LLMs are now primarily deployed as proprietary APIs Achiam et al. (2023); Team et al. (2023). Therefore, only prompt-based attacks are feasible, where adversaries have to operate through input manipulation. While existing attempts to compromise LLM reasoning Xu et al. (2024); Xiang et al. (2024) have demonstrated success in specific scenarios, they still face severe limitations in practice. A key challenge in attack design is to create attacks that are imperceptible to users. While obvious manipulations, such as altering final answers or inserting irrelevant steps, are easily detected by users, modifying the reasoning process while preserving narrative coherence is far more difficult. Existing methods often struggle to balance attack effectiveness with stealth, especially in the context of complex reasoning tasks.
Among the most relevant approaches, Xiang et al. (2024) employs misleading demonstrations to induce errors in LLMs. However, these methods are limited to in-context learning scenarios, requiring demonstrations as input, which limits their generalizability to the zero-shot settings. Furthermore, their strategy introduces an additional step that modifies the final answer, making it quite easy to identify by users. Another related approach, the preemptive answer “attack” Xu et al. (2024) alters the reasoning paradigm of the model by producing conclusions before deriving reasoning steps. Despite its novelty, this approach often generates easily identifiable outputs, reducing its imperceptibility and effectiveness in practice. These limitations are further evidenced by our experimental results in Section 3.2.
@@ 2.2 Overview of Stepwise Reasoning Error Disruption Attack
[…]
@@ 2.3 SEED-S: SEED Attack by Step Modification
Table 1: A comparison of the proportions of solutions generated by BadChain Xiang et al. (2024), UPA and MPA Xu et al. (2024), and SEED (SEED-S and SEED-P) that were detected by GPT-4o as originating from prompts containing attacks. The average improvement is determined by calculating the average decline in the detection rate of SEED compared to Xu et al. (2024).Z_S and F_S stands for the Zero-Shot and Few-Shot settings, respectively. Results demonstrate that SEED methods consistently achieve substantially lower detection rates across all model architectures and settings, with SEED-P showing particularly strong stealth capabilities while maintaining attack effectiveness.
@@
To further enhance the attack’s effectiveness, inspired by Xu et al. (2024), we prepend the corresponding incorrect answer to $R_{att}$. Finally, the modified output of the target LLM is obtained by feeding the modified problem’s incorrect answer and partial reasoning steps into it:
$q^{\prime}=LLM(D_{i}||I||p_{i}||a^{\prime}||R_{att}).$
Since we prepend $a^{\prime}$ to $R_{att}$, the proportion of $a^{\prime}$ relative to the entire input $q^{\prime}$ is minimal, and its position is central. Thus, we assume that its impact on the length of $R^{\prime}$ and the similarity between model outputs $R^{\prime}$ and $R[T_{att}:]$ is negligible.
@@ 3.1 Experimental Setup
Dataset. Building on prior studies targeting reasoning processes in LLMs Xu et al. (2024); Xiang et al. (2024), we evaluate our method using four datasets that encompass diverse and challenging reasoning tasks presented in two formats. Specifically, MATH Hendrycks et al. and GSM8K Cobbe et al. (2021) focus on arithmetic reasoning with open-ended formats, while MATHQA Amini et al. (2019) presents math problems in a multiple-choice format. CSQA Talmor et al. (2019), on the other hand, features multiple-choice commonsense reasoning tasks. As for the budget constraints, we follow the approach of Xiang et al. (2024), randomly sampling 500 questions from each dataset for our experiments. Further details about datasets are provided in Appendix A.
@@
Baselines. To our knowledge, UPA and MPA, introduced by Xu et al. (2024), along with BadChain Xiang et al. (2024), are the only methods targeting attacks on LLM reasoning. UPA and MPA prompt the LLM to generate an answer before the reasoning steps, with MPA further introducing a false answer to mislead reasoning. While BadChain achieves an ASR close to 100% across all datasets, its effectiveness is limited to the few-shot setting. Moreover, as Table 1 shows, its detection ratio nears 100% since it only modifies the final answer, warranting its exclusion from further discussion. Additionally, we find that the “Adding Mistake” method in Lanham et al. (2023) shares similarities with SEED-S, in that it introduces misleading reasoning steps. However, the “Adding Mistake” approach primarily focuses on examining whether CoT reasoning is post-hoc, rather than attack the reasoning of LLMs. Since the task of “Adding Mistake” differs from our single-round reasoning task, we concentrate solely on comparing the effectiveness of the attack.
Table 2: Comparison of performance measured by ASR under the setting in Xu et al. (2024). UPA and MPA are the methods proposed by Xu et al. (2024). Z_S and F_S stands for the Zero-Shot and Few-Shot settings, respectively. Highest ASR are highlighted within each model for a given dataset setting.
@@ 3.2 Overall Performance
As outlined in Section 1, we hypothesize that modifying $I_{solve}$ will lead to LLM outputs that are more easily identifiable by human, thereby diminishing the covert nature of the attack. To test this hypothesis, we evaluated the detection rates of SEED-generated solutions against baseline methods using GPT-4o to simulate human detection following the LLM-as-a-judge paradigm (Gu et al., 2024). Specifically, GPT-4o was employed to classify whether the input prompt had been attacked by analyzing the corresponding output solutions. Table 1 presents detection rates on the MATH and GSM8K datasets, demonstrating that UPA and MPA fail to satisfy the covert requirements. In contrast, both SEED-S and SEED-P exhibit substantial improvements in detection rates, particularly with GPT-4o, where the average improvement exceeds 90%. This validates that SEED achieves a higher level of stealth by preserving a natural reasoning flow without apparent manipulations.
@@ 3.3 Parameter Analysis
Conversely, higher $\sigma$ values also cause noticeable ASR drops in some cases, particularly with GPT-4o and Qwen-2.5 on MATH. We hypothesize that over-injecting reasoning steps can make the LLM more robust. When too many prior steps are introduced, the LLM focuses more on reviewing its prior reasoning rather than continuing with subsequent inference. This increased scrutiny helps the LLM detect inconsistencies and attempt corrections, leading to a more cautious reasoning approach and reducing the attack’s effectiveness. Additional results are provided in Appendix I due to space limitations.
@@ 4.1 Reasoning of LLMs
Enhancing reasoning in large language models (LLMs) remains a key research focus Yang et al. (2024); Ning et al. . The Chain of Thought (CoT) paradigm has been particularly effective, as shown by Wei et al. (2022) and Kojima et al. (2022), demonstrating that explicit reasoning steps, such as exemplars or step-by-step instructions, improve LLM performance. Subsequent work refined CoT with techniques like self-consistency Wang et al. , which uses majority voting across reasoning paths, and Least-to-Most Zhou et al. , a two-stage problem decomposition approach. Further extensions to trees Yao et al. (2024) and graphs Besta et al. (2024) expand CoT’s capabilities. Recent advances in long reasoning methods require LLMs to iteratively build upon prior steps, facilitating reflection Madaan et al. (2024); Zhao et al. (2024) or tree search Guan et al. (2025); Zhang et al. (2024) for subsequent reasoning steps, further expand the reasoning ability of LLMs. This reliance on step-by-step reasoning, however, raises new concerns regarding the vulnerability of LLMs.
@@ 4.2 Prompt-based Attack on LLMs
A key area of research aimed at ensuring the safety and robustness of LLMs involves developing methods to attack these models, prompting the generation of undesirable content Deng et al. (2023); Chu et al. (2024); Yu et al. (2024). One prominent category within this field focuses on "jailbreak" attacks, which bypass alignment mechanisms to elicit harmful or unsafe outputs Yi et al. (2024); Mehrotra et al. (2023); Zheng et al. (2024). However, our work is not directly related to jailbreak attacks. Instead, we focus on adversarial attacks, which subtly manipulate outputs without noticeable input modifications Xu et al. (2022); Kandpal et al. ; Xu et al. . While earlier studies targeted traditional NLP tasks such as sentiment analysis and classification Wang et al. (2024); Zhao et al. (2023), recent efforts have increasingly focused on attacking LLM reasoning processes Xiang et al. (2024); Xu et al. (2024). BadChain leverages backdoor vulnerabilities by embedding triggers within in-context learning demonstrations, but its applicability remains limited to specific contexts Xiang et al. (2024). Moreover, a critical drawback of BadChain is its nearly 100% detection rate, rendering it unsuitable for practical deployment. Similarly, UPA and MPA methods proposed by Xu et al. (2024), which instruct LLMs to generate answers before reasoning, often yield outputs that are easily identifiable, compromising their covert nature. Therefore, these approaches struggle to strike an effective balance between attack potency and stealth.
@@ Appendix A Details for the Datasets
Figure 6: The form used for human evaluation of coverage of each attack approach.
@@ Appendix B Implementation of SEED attack
In Figure 4, we present the prompts employed for both the attack and problem-solving across different tasks. Additionally, in Figure 10, we display the demonstrations used in Few-Shot settings for each dataset. For a fair evaluation, it is important to note that we utilized the same demonstrations as those in Xu et al. (2024).
@@ Appendix H Ablation Study
Appendix I More Experiment Results

### v5

# Stepwise Reasoning Disruption Attack of LLMs

@@ Abstract
Large language models (LLMs) have made remarkable strides in complex reasoning tasks, but their safety and robustness in reasoning processes remain unexplored, particularly in third-party platforms that facilitate user interactions via APIs. Existing attacks on LLM reasoning are constrained by specific settings or lack of imperceptibility, limiting their feasibility and generalizability. To address these challenges, we propose the Stepwise rEasoning Error Disruption (SEED) attack, which subtly injects errors into prior reasoning steps to mislead the model into producing incorrect subsequent reasoning and final answers. Unlike previous methods, SEED is compatible with zero-shot and few-shot settings, maintains the natural reasoning flow, and ensures covert execution without modifying the instruction. Extensive experiments on four datasets across four different models demonstrate SEED’s effectiveness, revealing the vulnerabilities of LLMs to disruptions in reasoning processes. These findings underscore the need for greater attention to the robustness of LLM reasoning to ensure safety in practical applications. Our code is available at: https://github.com/Applied-Machine-Learning-Lab/SEED-Attack
@@
Stepwise Reasoning Disruption Attack of LLMs
Jingyu Peng‡§111 Both authors contributed equally to this research., Maolin Wang§111 Both authors contributed equally to this research., Xiangyu Zhao§ 222 Corresponding authors. , Kai Zhang, Wanyu Wang§, Pengyue Jia§, Qidong Liu§, Ruocheng Guo, Qi Liu 222 Corresponding authors. University of Science and Technology of China, § City University of Hong Kong, Independent Researcher jypeng28@mail.ustc.edu.cn, xianzhao@cityu.edu.hk {morin.wang, jia.pengyue, wanyuwang4-c}@my.cityu.edu.hk {kkzhang08, qiliuql}@ustc.edu.cn, rguo.asu@gmail.com, liuqidong@stu.xjtu.edu.cn
@@ 1 Introduction
Large language models (LLMs) have remarkably improved complex tasks by adopting various enhanced reasoning approaches Besta et al. (2024); Xu et al. (2024b); Yang et al. (2024); Yao et al. (2024); Xu et al. (2024a); Jia et al. (2024); Cheng et al. (2024). These approaches have boosted their performance and drawn attention to the trustworthiness of the reasoning processes, including faithfulness Lanham et al. (2023); Turpin et al. (2024), fairness Shaikh et al. (2023), and safety Xu et al. (2024c).
@@
Exploiting such vulnerability in LLMs introduces two fundamental challenges: feasibility and imperceptibility. Technically, unlike traditional adversarial attack methods, which often leverage internal information of target models such as gradients and logits, state-of-the-art LLMs are now primarily deployed as proprietary APIs Achiam et al. (2023); Team et al. (2023). Therefore, only prompt-based attacks are feasible, where adversaries have to operate through input manipulation. While existing attempts to compromise LLM reasoning Xu et al. (2024c); Xiang et al. (2024); Ni et al. (2025) have demonstrated success in specific scenarios, they still face severe limitations in practice. A key challenge in attack design is to create attacks that are imperceptible to users. While obvious manipulations, such as altering final answers or inserting irrelevant steps, are easily detected by users, modifying the reasoning process while preserving narrative coherence is far more difficult. Existing methods often struggle to balance attack effectiveness with stealth, especially in the context of complex reasoning tasks.
Among the most relevant approaches, Xiang et al. (2024) employs misleading demonstrations to induce errors in LLMs. However, these methods are limited to in-context learning scenarios, requiring demonstrations as input, which limits their generalizability to the zero-shot settings. Furthermore, their strategy introduces an additional step that modifies the final answer, making it quite easy to identify by users. Another related approach, the preemptive answer “attack” Xu et al. (2024c) alters the reasoning paradigm of the model by producing conclusions before deriving reasoning steps. Despite its novelty, this approach often generates easily identifiable outputs, reducing its imperceptibility and effectiveness in practice. These limitations are further evidenced by our experimental results in Section 3.2.
@@ 2.2 Overview of Stepwise Reasoning Error Disruption Attack
Therefore, our work focuses on how to implement a $M(\cdot)$ where $R_{att}=M(p)$, that satisfy the variation of Eq. 1:
@@ 2.3 SEED-S: SEED Attack by Step Modification
Table 1: A comparison of the proportions of solutions generated by BadChain Xiang et al. (2024), UPA and MPA Xu et al. (2024c), and SEED (SEED-S and SEED-P) that were detected by GPT-4o as originating from prompts containing attacks. The average improvement is determined by calculating the average decline in the detection rate of SEED compared to Xu et al. (2024c).Z_S and F_S stands for the Zero-Shot and Few-Shot settings, respectively. Results demonstrate that SEED methods consistently achieve substantially lower detection rates across all model architectures and settings, with SEED-P showing particularly strong stealth capabilities while maintaining attack effectiveness.
@@
To further enhance the attack’s effectiveness, inspired by Xu et al. (2024c), we prepend the corresponding incorrect answer $a_{mod}$ to $R_{att}$. Finally, the modified output of the target LLM is obtained by feeding the modified problem’s incorrect answer and partial reasoning steps into it:
$q^{\prime}=LLM(I||D||p||a_{mod}||R_{att}).$
Since we prepend $a_{mod}$ to $R_{att}$, the proportion of $a_{mod}$ relative to the entire input $q^{\prime}$ is minimal, and its position is central. Thus, we assume that its impact on the length of $R^{\prime}$ and the similarity between model outputs $R^{\prime}$ and $R[T_{att}:]$ is negligible.
@@ 3.1 Experimental Setup
Dataset. Building on prior studies targeting reasoning processes in LLMs Xu et al. (2024c); Xiang et al. (2024), we evaluate our method using four datasets that encompass diverse and challenging reasoning tasks presented in two formats. Specifically, MATH Hendrycks et al. and GSM8K Cobbe et al. (2021) focus on arithmetic reasoning with open-ended formats, while MATHQA Amini et al. (2019) presents math problems in a multiple-choice format. CSQA Talmor et al. (2019), on the other hand, features multiple-choice commonsense reasoning tasks. As for the budget constraints, we follow the approach of Xiang et al. (2024), randomly sampling 500 questions from each dataset for our experiments. Further details about datasets are provided in Appendix A.
@@
Baselines. To our knowledge, UPA and MPA, introduced by Xu et al. (2024c), along with BadChain Xiang et al. (2024), are the only methods targeting attacks on LLM reasoning. UPA and MPA prompt the LLM to generate an answer before the reasoning steps, with MPA further introducing a false answer to mislead reasoning. While BadChain achieves an ASR close to 100% across all datasets, its effectiveness is limited to the few-shot setting. Moreover, as Table 1 shows, its detection ratio nears 100% since it only modifies the final answer, warranting its exclusion from further discussion. Additionally, we find that the “Adding Mistake” method in Lanham et al. (2023) shares similarities with SEED-S, in that it introduces misleading reasoning steps. However, the “Adding Mistake” approach primarily focuses on examining whether CoT reasoning is post-hoc, rather than attack the reasoning of LLMs. Since the task of “Adding Mistake” differs from our single-round reasoning task, we concentrate solely on comparing the effectiveness of the attack.
Table 2: Comparison of performance measured by ASR under the setting in Xu et al. (2024c). UPA and MPA are the methods proposed by Xu et al. (2024c). Z_S and F_S stands for the Zero-Shot and Few-Shot settings, respectively. Highest ASR are highlighted within each model for a given dataset setting.
@@ 3.2 Overall Performance
As outlined in Section 1, we hypothesize that modifying $I_{solve}$ will lead to LLM outputs that are more easily identifiable by human, thereby diminishing the covert nature of the attack. To test this hypothesis, we evaluated the detection rates of SEED-generated solutions against baseline methods using GPT-4o to simulate human detection following the LLM-as-a-judge paradigm (Gu et al., 2024). Specifically, GPT-4o was employed to classify whether the input prompt had been attacked by analyzing the corresponding output solutions. Table 1 presents detection rates on the MATH and GSM8K datasets, demonstrating that UPA and MPA fail to satisfy the covert requirements. In contrast, both SEED-S and SEED-P exhibit substantial improvements in detection rates, particularly with GPT-4o, where the average improvement exceeds 90%. This validates that SEED achieves a higher level of stealth by preserving a natural reasoning flow without apparent manipulations. Meanwhile, SEED-S achieves lower detection rates than SEED-P, indicating superior stealth capabilities, as SEED-S introduces less modification to the prior reasoning process.
@@
We also assessed the detection rate for successfully attacked solutions, with the results presented in Section I.
@@ 3.3 Parameter Analysis
Conversely, higher $\sigma$ values also cause noticeable ASR drops in some cases, particularly with GPT-4o and Qwen-2.5 on MATH. We hypothesize that over-injecting reasoning steps can make the LLM more robust. When too many prior steps are introduced, the LLM focuses more on reviewing its prior reasoning rather than continuing with subsequent inference. This increased scrutiny helps the LLM detect inconsistencies and attempt corrections, leading to a more cautious reasoning approach and reducing the attack’s effectiveness. Additional results are provided in Appendix J due to space limitations.
@@ 4.1 Reasoning of LLMs
Enhancing reasoning in large language models (LLMs) remains a key research focus Yang et al. (2024); Ning et al. ; Li et al. (2023); Yuan et al. (2024); Liu et al. (2024). The Chain of Thought (CoT) paradigm has been particularly effective, as shown by Wei et al. (2022) and Kojima et al. (2022), demonstrating that explicit reasoning steps, such as exemplars or step-by-step instructions, improve LLM performance. Subsequent work refined CoT with techniques like self-consistency Wang et al. , which uses majority voting across reasoning paths, and Least-to-Most Zhou et al. , a two-stage problem decomposition approach. Further extensions to trees Yao et al. (2024) and graphs Besta et al. (2024) expand CoT’s capabilities. Recent advances in long reasoning methods require LLMs to iteratively build upon prior steps, facilitating reflection Madaan et al. (2024); Zhao et al. (2024) or tree search Guan et al. (2025); Zhang et al. (2024, 2022) for subsequent reasoning steps, further expand the reasoning ability of LLMs. This reliance on step-by-step reasoning, however, raises new concerns regarding the vulnerability of LLMs.
@@ 4.2 Prompt-based Attack on LLMs
A key area of trustworthy AI research Fan et al. (2023); Chen et al. (2022); Jia et al. (2024); Xu et al. (2025) aimed at ensuring the safety and robustness of LLMs involves developing methods to attack these models, prompting the generation of undesirable content Deng et al. (2023); Chu et al. (2024); Yu et al. (2024); Zhang et al. (2021). One prominent category within this field focuses on "jailbreak" attacks, which bypass alignment mechanisms to elicit harmful or unsafe outputs Yi et al. (2024); Mehrotra et al. (2023); Zheng et al. (2024). However, our work is not directly related to jailbreak attacks. Instead, we focus on adversarial attacks, which subtly manipulate outputs without noticeable input modifications Xu et al. (2022); Kandpal et al. ; Xu et al. . While earlier studies targeted traditional NLP tasks such as sentiment analysis and classification Wang et al. (2024); Zhao et al. (2023); Zhang et al. (2019), recent efforts have increasingly focused on attacking LLM reasoning processes Xiang et al. (2024); Xu et al. (2024c). BadChain leverages backdoor vulnerabilities by embedding triggers within in-context learning demonstrations, but its applicability remains limited to specific contexts Xiang et al. (2024). Moreover, a critical drawback of BadChain is its nearly 100% detection rate, rendering it unsuitable for practical deployment. Similarly, UPA and MPA methods proposed by Xu et al. (2024c), which instruct LLMs to generate answers before reasoning, often yield outputs that are easily identifiable, compromising their covert nature. Therefore, these approaches struggle to strike an effective balance between attack potency and stealth.
@@ 6 Limitation
7 Acknowledge
This research was partially supported by the National Natural Science Foundation of China (Grants No.62406303), Anhui Province Science and Technology Innovation Project (202423k09020010), Research Impact Fund (No.R1015-23), Collaborative Research Fund (No.C1043-24GF),the Fundamental Research Funds for the Central Universities, Huawei (Huawei Innovation Research Program, Huawei Fellowship), Tencent (CCF-Tencent Open Fund, Tencent Rhino-Bird Focused Research Program), Alibaba (CCF-Alimama Tech Kangaroo Fund No. 2024002), Ant Group (CCF-Ant Research Fund), and Kuaishou.
@@ Appendix A Details for the Datasets
Figure 6: The form used for human evaluation of covert of each attack approach.
@@ Appendix B Implementation of SEED attack
In Figure 4, we present the prompts employed for both the attack and problem-solving across different tasks. Additionally, in Figure 10, we display the demonstrations used in Few-Shot settings for each dataset. For a fair evaluation, it is important to note that we utilized the same demonstrations as those in Xu et al. (2024c).
@@ Appendix H Ablation Study
Table 7: Comparison of the detection rate in successfully attacked solutions.
MATH GSM8K
Qwen BadChain 1.000 1.000
UPA 0.600 0.628
MPA 0.741 0.724
SEED-S 0.098 0.131
SEED-P 0.216 0.344
GPT BadChain 1.000 1.000
UPA 0.858 0.86
MPA 0.756 0.844
SEED-S 0.045 0.012
SEED-P 0.148 0.144
Appendix I The comparison of detection rate on successfully attacked solutions
In Table 7, we provide a comparison of the detection rate in successfully attacked solutions. The average detection rate is significantly higher than the overall attack detection rate, which aligns with the intuition that only modified solutions have the potential to be detected (Since ASR is computed based on questions where the original answer was correct, the detection rate for failed attacks is not zero). Additionally, the comparison between SEED and the baselines is consistent with the results presented in Table 1.
Appendix J More Experiment Results